# %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

### v6

@@ Abstract
This paper presents a self-supervised temporal video alignment framework which is useful for several fine-grained human activity understanding applications. In contrast with the state-of-the-art method of CASA, where sequences of 3D skeleton coordinates are taken directly as input, our key idea is to use sequences of 2D skeleton heatmaps as input. Unlike CASA which performs self-attention in the temporal domain only, we feed 2D skeleton heatmaps to a video transformer which performs self-attention both in the spatial and temporal domains for extracting effective spatiotemporal and contextual features. In addition, we introduce simple heatmap augmentation techniques based on 2D skeletons for self-supervised learning. Despite the lack of 3D information, our approach achieves not only higher accuracy but also better robustness against missing and noisy keypoints than CASA. Furthermore, extensive evaluations on three public datasets, i.e., Penn Action, IKEA ASM, and H2O, demonstrate that our approach outperforms previous methods in different fine-grained human activity understanding tasks. Finally, fusing 2D skeleton heatmaps with RGB videos yields the state-of-the-art on all metrics and datasets. To our best knowledge, our work is the first to utilize 2D skeleton heatmap inputs and the first to explore multi-modality fusion for temporal video alignment.
Keywords:
Temporal video alignment Temporal 2D skeleton sequence alignment Multi-modality fusion Self-supervised learning
@@ 1 Introduction
We study the problem of temporal video alignment [62, 23, 30, 50, 43], which aims to find dense framewise correspondences between two videos capturing the same activity, despite discrepancies in human appearance, background clutter, human motion, and camera viewpoint. Temporal video alignment plays an important role in many fine-grained human activity understanding applications, including anomaly detection [23] (i.e., detect deviations from a dataset of normal videos) and skill transfer [43] (i.e., provide feedback to learners based on expert demonstration videos).
We are interested in self-supervised approaches for temporal video alignment, since acquiring fine-grained correspondence labels is generally hard and costly [23, 38]. Early methods [62, 23, 30, 50] attempt to match video frames directly. They optimize different temporal alignment metrics such as dynamic time warping [18] and optimal transport [17] for self-supervised learning. Although the learned representations show promising generalization to downstream tasks, the results are far from being practical. Recently, CASA [43] proposes a context-aware self-supervised method for temporal video alignment, which operates on 3D skeletons (see Fig. 1(a)). It conducts 3D skeleton-based augmentations for self-supervised learning, while employing a transformer [72] to learn contextual features, yielding great results. Since CASA [43] directly processes 3D skeleton coordinates, its performance is greatly affected by human pose estimation errors such as noisy and missing keypoints.
We present a self-supervised temporal video alignment method, which relies on 2D skeletons (see Fig. 1(b)). Motivated by the recent successes of [22, 34], we convert sequences of 2D skeletons into sequences of heatmaps, where each joint is modeled as a Gaussian distribution, making it less sensitive to human pose estimation errors. Unlike CASA [43] which performs self-attention in the temporal domain only, we pass the heatmap sequences to our alignment network, which adopts a video transformer [2] to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual features. Moreover, we propose simple various heatmap augmentation strategies based on 2D skeletons for self-supervised learning. As compared to CASA [43], our method based on 2D skeleton heatmaps obtains both higher accuracy and better robustness against missing and noisy keypoints despite lacking 3D information. Moreover, we show superior performance of our approach over previous methods in fine-grained human activity understanding tasks on three public datasets. Finally, utilizing both 2D skeleton heatmaps and RGB videos as inputs achieves the state-of-the-art across all metrics and datasets.
@@
  • Our 2D skeleton-based approach is more accurate and robust against noisy and missing keypoints than CASA [43] which requires 3D skeletons. Our evaluations on Penn Action, IKEA ASM, and H2O show that our approach outperforms previous methods on fine-grained human activity understanding tasks.
@@ 2 Related Work
Self-Supervised Learning. Considerable efforts have been made in developing pretext tasks with pseudo labels for training image-based self-supervised models. Examples include image colorization [44, 45], object counting [57, 51], predicting rotations [26], solving puzzles [9, 39], image inpainting [35], and image clustering [10, 11]. Recently, several works [12, 13, 28, 14] focus on designing data augmentation and contrastive learning strategies, which are important for learning effective representations. Unlike the above image-based methods, video-based methods, which exploit both spatial cues and temporal cues in videos, have attracted a great amount of interests recently. They learn representations by using pretext tasks such as future frame prediction [67, 73, 1, 21] and frame clustering [41, 71], or leveraging temporal information such as temporal coherence [56, 84, 27], temporal order [54, 46, 25, 76], arrow of time [59, 75], and pace [5, 74, 79], or utilizing contrastive learning techniques [24, 33, 61, 20].
@@
More recently, skeleton-based methods have been introduced. They learn representations by using pretext tasks such as skeleton inpainting [82] and motion prediction [68], or utilizing neighborhood consistency [64], motion continuity [69], and multiple pretext tasks [48]. The above methods often do not exploit spatiotemporal dependencies. Here, we use 2D skeleton-based alignment as a pretext task and 2D skeleton-based augmentation for self-supervised learning.
@@
Temporal Video Alignment. Despite significant progress in unsupervised time series alignment, methods for self-supervised temporal video alignment have emerged only recently. To learn representations, TCN [62] employs contrastive learning on synchronized frames across viewpoints, while TCC [23] imposes cycle-consistent frame correspondences across videos. Recently, LAV [30] and VAVA [50] adopt dynamic time warping [18] and optimal transport [17] respectively as temporal video alignment losses. The above methods focus on aligning video frames directly. More recently, CASA [43] presents a self-supervised temporal video alignment method which operates on 3D skeletons. In this work, we propose a 2D skeleton-based self-supervised temporal video alignment method, which is more accurate and robust against noisy and missing keypoints. We further fuse 2D skeleton heatmaps with RGB videos, establishing the state-of-the-art performance on all metrics and datasets.
@@
Skeleton-Based Action Recognition. There exist notable interests in skeleton-based action recognition due to its compactness. A popular group of methods [78, 65, 7, 15, 29, 66] model the skeleton sequence as a spatiotemporal graph and rely on Graph Convolutional Networks (GCNs). Despite great results, their drawbacks lie in robustness to noisy/incomplete skeletons [83], scalability with multi-person actions, and fusion with other modalities [19]. Another group of methods convert the skeleton sequence into a 2D input (e.g., heatmap/coordinate aggregation [37, 16, 47, 52, 77, 6, 3]) or a 3D input (e.g., stacked pseudo images [32, 49]) and employ Convolutional Neural Networks (CNNs). These methods often suffer from information loss during aggregation. To address that, Duan et al. [22] convert each skeleton into a heatmap and pass the heatmap sequence, which preserves all information, to a CNN, and achieve superior results while alleviating the above issues of GCN-based methods. Inspired by their method for coarse-grained action recognition, we propose a self-supervised temporal video alignment framework for fine-grained tasks, e.g., action phase classification. In addition to video-level cues, fine-grained tasks further require frame-level cues and architectures/losses for extracting/utilizing them, which makes them more challenging.
@@ 3 Our Approach
We discuss the estimation of 2D skeletons $\bm{S}$ and their corresponding heatmaps $\bm{H}$ from a video sequence. An off-the-shelf 2D human pose estimator such as OpenPose [8] can be used to extract good quality 2D skeletons. For memory efficiency, a 2D skeleton is often stored as a set of 2D joint triplets $\{(x_{k},y_{k},c_{k})\}$, where $(x_{k},y_{k})$ are the 2D coordinates and $c_{k}$ is the (maximum) confidence score of the $k$-th joint. In practice, for fair comparisons, we use 2D skeletons that are projected (via orthographic projection) from 3D skeletons provided by CASA [43] and set $c_{k}=1.0$. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). Our results with OpenPose 2D poses can be found in the supplementary material.
@@
where the $l$-th limb is the segment $seg(a_{l},b_{l})$ between the two joints $a_{l}$ and $b_{l}$, and the $dist$ function computes the distance from the location $(i,j)$ to the segment $seg(a_{l},b_{l})$. As we will show later in Sec. 4.1.2, combining both joint and limb heatmaps as input to our alignment network yields the best results, i.e., for each 2D skeleton $\bm{s}_{i}$, we extract the combined heatmap $\bm{h}_{i}=\bm{h}^{J+L}_{i}$. Moreover, since the action of interest appears only in a small region in the video frames, we crop the heatmaps along the spatial dimensions using the smallest bounding box containing all 2D skeletons in the video sequence and resize them to the spatial dimensions of $30\times 30$. Thus, we reduce the size of the heatmaps, while preserving all 2D skeletons and their motions.
2D Skeleton Heatmaps vs. 3D Skeleton Coordinates. Unlike CASA [43] which operates on 3D skeleton coordinates, our method relies on 2D skeleton heatmaps. Advantages of 2D skeleton heatmaps over 3D skeleton coordinates include: 1) Accurate 2D skeletons are generally easier to obtain, as accurate depths are not required (e.g., see Fig. 1 of [22]), 2) 2D skeleton heatmaps are more robust, as joints/limbs are modeled by Gaussian maps, 3) 2D skeleton heatmaps are applicable to advanced CNNs, whereas MLPs/GCNs are often used for 3D skeleton coordinates, and 4) It is convenient to fuse 2D skeleton heatmaps with other grid-like data. The above advantages contribute to our superior accuracy and robustness, as shown later in Sec. 4.
@@
Below we describe our simple heatmap augmentation techniques based on 2D skeletons, which are used to generate correspondence labels for self-supervised training of our alignment network. This is in contrast with the coordinate augmentation techniques in CASA [43], which rely on 3D skeletons. We first apply various augmentations on the original 2D skeletons $\bm{S}$ to obtain the augmented 2D skeletons $\bm{S}^{\prime}$, and then compute the augmented heatmaps $\bm{H}^{\prime}$ from the augmented 2D skeletons $\bm{S}^{\prime}$. They include:
@@
Fig. 3 shows examples of our simple heatmap augmentation techniques based on 2D skeletons. For position and orientation augmentations, we empirically use a multivariate Gaussian distribution with a covariance matrix so that the added noise is temporally smooth as in CASA [43]. However, it does not lead to improved results and is thus not used in the above. Furthermore, the above augmentations generate correspondence labels between $\bm{S}$ and $\bm{S}^{\prime}$ (and similarly, between $\bm{H}$ and $\bm{H}^{\prime}$), which can be used for self-supervised training of our alignment network in the next section.
@@
Self-Attention. Given the original heatmap sequence $\bm{H}$ and the augmented sequence $\bm{H}^{\prime}$, we pass them separately through a video transformer, namely ViViT [2], to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual cues within each sequence. Below we discuss the processing of $\bm{H}$ only since the same is applied for $\bm{H}^{\prime}$. For each $\bm{h}_{i}$ in $\bm{H}$, we first break it into $h*w$ non-overlapping patches, which are embedded via linear projection into $h*w$ tokens. The tokens from $\bm{h}_{i}$ (with spatial positional encoding) are then fed to a spatial transformer, which learns interactions between tokens from the same temporal index (i.e., between joints in the same skeleton). Next, we employ global average pooling on the above tokens output by the spatial transformer to obtain the frame-level token representing $\bm{h}_{i}$. The frame-level tokens of $\bm{H}$ (with temporal positional encoding) are passed together to a temporal transformer, which learns interactions between tokens from different temporal indexes (i.e., between skeletons in the same sequence). Note that our approach performs self-attention both in the spatial and temporal domains, whereas CASA [43] conducts self-attention in the temporal domain only. Fig. 4 illustrates this difference. In addition, our positional encodings are learned, whereas it is fixed in CASA [43].
Cross-Attention. The above modules only focus on self-attention in the spatial and temporal domains within each sequence. Following [70, 43], we add a cross-attention module to learn interactions between $\bm{H}$ and $\bm{H}^{\prime}$ (i.e., between skeletons across the sequences), yielding contextual cues across the sequences. In particular, we encode the tokens output by the self-attention modules via an MLP layer and positional encoding before applying cross-attention.
Projection Head. We include a projection head, which is an MLP network with one hidden layer. As mentioned in [13, 43], adding the projection head improves the generalization ability and yields effective features for downstream tasks.
Matching and Loss. We first compute the probability $\bm{\omega}_{ij}$ that the $i$-th frame in $\bm{H}$ is matched with the $j$-th frame in $\bm{H}^{\prime}$ using contrastive regression [23, 43] as:
$\bm{\omega}_{ij}=\frac{e^{-||\bm{z}_{i}-\bm{z}^{\prime}_{j}||/\lambda}}{\sum^{M}_{m=1}e^{-||\bm{z}_{m}-\bm{z}^{\prime}_{j}||/\lambda}},$ (3)

### v7

@@ Abstract
This paper presents a self-supervised temporal video alignment framework which is useful for several fine-grained human activity understanding applications. In contrast with the state-of-the-art method of CASA, where sequences of 3D skeleton coordinates are taken directly as input, our key idea is to use sequences of 2D skeleton heatmaps as input. Unlike CASA which performs self-attention in the temporal domain only, we feed 2D skeleton heatmaps to a video transformer which performs self-attention both in the spatial and temporal domains for extracting effective spatiotemporal and contextual features. In addition, we introduce simple heatmap augmentation techniques based on 2D skeletons for self-supervised learning. Despite the lack of 3D information, our approach achieves not only higher accuracy but also better robustness against missing and noisy keypoints than CASA. Furthermore, extensive evaluations on three public datasets, i.e., Penn Action, IKEA ASM, and H2O, demonstrate that our approach outperforms previous methods in different fine-grained human activity understanding tasks. Finally, fusing 2D skeleton heatmaps with RGB videos yields the state-of-the-art on all metrics and datasets. To our best knowledge, our work is the first to utilize 2D skeleton heatmap inputs and the first to explore multi-modality fusion for temporal video alignment. Our code and dataset are available on our research website: https://retrocausal.ai/research/.
@@ 1 Introduction
We study the problem of temporal video alignment [sermanet2018tcn, dwibedi2019tcc, haresh2021learning, liu2022learning, kwon2022context], which aims to find dense framewise correspondences between two videos capturing the same activity, despite discrepancies in human appearance, background clutter, human motion, and camera viewpoint. Temporal video alignment plays an important role in many fine-grained human activity understanding applications, including anomaly detection [dwibedi2019tcc] (i.e., detect deviations from a dataset of normal videos) and skill transfer [kwon2022context] (i.e., provide feedback to learners based on expert demonstration videos).
We are interested in self-supervised approaches for temporal video alignment, since acquiring fine-grained correspondence labels is generally hard and costly [dwibedi2019tcc, khan2022timestamp]. Early methods [sermanet2018tcn, dwibedi2019tcc, haresh2021learning, liu2022learning] attempt to match video frames directly. They optimize different temporal alignment metrics such as dynamic time warping [cuturi2017soft] and optimal transport [cuturi2013sinkhorn] for self-supervised learning. Although the learned representations show promising generalization to downstream tasks, the results are far from being practical. Recently, CASA [kwon2022context] proposes a context-aware self-supervised method for temporal video alignment, which operates on 3D skeletons (see Fig. 1(a)). It conducts 3D skeleton-based augmentations for self-supervised learning, while employing a transformer [vaswani2017attention] to learn contextual features, yielding great results. Since CASA [kwon2022context] directly processes 3D skeleton coordinates, its performance is greatly affected by human pose estimation errors such as noisy and missing keypoints.
We present a self-supervised temporal video alignment method, which relies on 2D skeletons (see Fig. 1(b)). Motivated by the recent successes of [duan2022revisiting, hyder2024action], we convert sequences of 2D skeletons into sequences of heatmaps, where each joint is modeled as a Gaussian distribution, making it less sensitive to human pose estimation errors. Unlike CASA [kwon2022context] which performs self-attention in the temporal domain only, we pass the heatmap sequences to our alignment network, which adopts a video transformer [arnab2021vivit] to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual features. Moreover, we propose simple various heatmap augmentation strategies based on 2D skeletons for self-supervised learning. As compared to CASA [kwon2022context], our method based on 2D skeleton heatmaps obtains both higher accuracy and better robustness against missing and noisy keypoints despite lacking 3D information. Moreover, we show superior performance of our approach over previous methods in fine-grained human activity understanding tasks on three public datasets. Finally, utilizing both 2D skeleton heatmaps and RGB videos as inputs achieves the state-of-the-art across all metrics and datasets.
@@
  • Our 2D skeleton-based approach is more accurate and robust against noisy and missing keypoints than CASA [kwon2022context] which requires 3D skeletons. Our evaluations on Penn Action, IKEA ASM, and H2O show that our approach outperforms previous methods on fine-grained human activity understanding tasks.
@@ 2 Related Work
Self-Supervised Learning. Considerable efforts have been made in developing pretext tasks with pseudo labels for training image-based self-supervised models. Examples include image colorization [larsson2016learning, larsson2017colorization], object counting [noroozi2017representation, liu2018leveraging], predicting rotations [gidaris2018unsupervised], solving puzzles [carlucci2019domain, kim2019self], image inpainting [jenni2020steering], and image clustering [caron2018deep, caron2019unsupervised]. Recently, several works [caron2020unsupervised, chen2020simple, grill2020bootstrap, chen2021exploring] focus on designing data augmentation and contrastive learning strategies, which are important for learning effective representations. Unlike the above image-based methods, video-based methods, which exploit both spatial cues and temporal cues in videos, have attracted a great amount of interests recently. They learn representations by using pretext tasks such as future frame prediction [srivastava2015unsupervised, vondrick2016generating, ahsan2018discrimnet, diba2019dynamonet] and frame clustering [kumar2022unsupervised, tran2024permutation], or leveraging temporal information such as temporal coherence [mobahi2009deep, zou2011unsupervised, goroshin2015unsupervised], temporal order [misra2016shuffle, lee2017unsupervised, fernando2017self, xu2019self], arrow of time [pickup2014seeing, wei2018learning], and pace [benaim2020speednet, wang2020self, yao2020video], or utilizing contrastive learning techniques [feichtenhofer2021large, hu2021contrast, qian2021spatiotemporal, dave2022tclr].
@@
More recently, skeleton-based methods have been introduced. They learn representations by using pretext tasks such as skeleton inpainting [zheng2018unsupervised] and motion prediction [su2020predict], or utilizing neighborhood consistency [si2020adversarial], motion continuity [su2021self], and multiple pretext tasks [lin2020ms2l]. The above methods often do not exploit spatiotemporal dependencies. Here, we use 2D skeleton-based alignment as a pretext task and 2D skeleton-based augmentation for self-supervised learning.
@@
Temporal Video Alignment. Despite significant progress in unsupervised time series alignment, methods for self-supervised temporal video alignment have emerged only recently. To learn representations, TCN [sermanet2018tcn] employs contrastive learning on synchronized frames across viewpoints, while TCC [dwibedi2019tcc] imposes cycle-consistent frame correspondences across videos. Recently, LAV [haresh2021learning] and VAVA [liu2022learning] adopt dynamic time warping [cuturi2017soft] and optimal transport [cuturi2013sinkhorn] respectively as temporal video alignment losses. The above methods focus on aligning video frames directly. More recently, CASA [kwon2022context] presents a self-supervised temporal video alignment method which operates on 3D skeletons. In this work, we propose a 2D skeleton-based self-supervised temporal video alignment method, which is more accurate and robust against noisy and missing keypoints. We further fuse 2D skeleton heatmaps with RGB videos, establishing the state-of-the-art performance on all metrics and datasets.
@@
Skeleton-Based Action Recognition. There exist notable interests in skeleton-based action recognition due to its compactness. A popular group of methods [yan2018spatial, song2020stronger, cai2021jolo, chen2021channel, gupta2021quo, song2022constructing] model the skeleton sequence as a spatiotemporal graph and rely on Graph Convolutional Networks (GCNs). Despite great results, their drawbacks lie in robustness to noisy/incomplete skeletons [zhu2019robust], scalability with multi-person actions, and fusion with other modalities [das2020vpn]. Another group of methods convert the skeleton sequence into a 2D input (e.g., heatmap/coordinate aggregation [ke2017new, choutas2018potion, li2018co, luvizon20182d, yan2019pa3d, caetano2019skelemotion, asghari2020dynamic]) or a 3D input (e.g., stacked pseudo images [hernandez20173d, lin2020image]) and employ Convolutional Neural Networks (CNNs). These methods often suffer from information loss during aggregation. To address that, Duan et al. [duan2022revisiting] convert each skeleton into a heatmap and pass the heatmap sequence, which preserves all information, to a CNN, and achieve superior results while alleviating the above issues of GCN-based methods. Inspired by their method for coarse-grained action recognition, we propose a self-supervised temporal video alignment framework for fine-grained tasks, e.g., action phase classification. In addition to video-level cues, fine-grained tasks further require frame-level cues and architectures/losses for extracting/utilizing them, which makes them more challenging.
@@ 3 Our Approach
We discuss the estimation of 2D skeletons $\bm{S}$ and their corresponding heatmaps $\bm{H}$ from a video sequence. An off-the-shelf 2D human pose estimator such as OpenPose [cao2019openpose] can be used to extract good quality 2D skeletons. For memory efficiency, a 2D skeleton is often stored as a set of 2D joint triplets $\{(x_{k},y_{k},c_{k})\}$, where $(x_{k},y_{k})$ are the 2D coordinates and $c_{k}$ is the (maximum) confidence score of the $k$-th joint. In practice, for fair comparisons, we use 2D skeletons that are projected (via orthographic projection) from 3D skeletons provided by CASA [kwon2022context] and set $c_{k}=1.0$. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). Our results with OpenPose 2D poses can be found in the supplementary material.
@@
where the $l$-th limb is the segment $seg(a_{l},b_{l})$ between the two joints $a_{l}$ and $b_{l}$, and the distdist function computes the distance from the location $(i,j)$ to the segment $seg(a_{l},b_{l})$. As we will show later in Sec. 4.1.2, combining both joint and limb heatmaps as input to our alignment network yields the best results, i.e., for each 2D skeleton $\bm{s}_{i}$, we extract the combined heatmap $\bm{h}_{i}=\bm{h}^{J+L}_{i}$. Moreover, since the action of interest appears only in a small region in the video frames, we crop the heatmaps along the spatial dimensions using the smallest bounding box containing all 2D skeletons in the video sequence and resize them to the spatial dimensions of $30\times 30$. Thus, we reduce the size of the heatmaps, while preserving all 2D skeletons and their motions.
2D Skeleton Heatmaps vs. 3D Skeleton Coordinates. Unlike CASA [kwon2022context] which operates on 3D skeleton coordinates, our method relies on 2D skeleton heatmaps. Advantages of 2D skeleton heatmaps over 3D skeleton coordinates include: 1) Accurate 2D skeletons are generally easier to obtain, as accurate depths are not required (e.g., see Fig. 1 of [22]), 2) 2D skeleton heatmaps are more robust, as joints/limbs are modeled by Gaussian maps, 3) 2D skeleton heatmaps are applicable to advanced CNNs, whereas MLPs/GCNs are often used for 3D skeleton coordinates, and 4) It is convenient to fuse 2D skeleton heatmaps with other grid-like data. The above advantages contribute to our superior accuracy and robustness, as shown later in Sec. 4.
@@
Below we describe our simple heatmap augmentation techniques based on 2D skeletons, which are used to generate correspondence labels for self-supervised training of our alignment network. This is in contrast with the coordinate augmentation techniques in CASA [kwon2022context], which rely on 3D skeletons. We first apply various augmentations on the original 2D skeletons $\bm{S}$ to obtain the augmented 2D skeletons $\bm{S}^{\prime}$, and then compute the augmented heatmaps $\bm{H}^{\prime}$ from the augmented 2D skeletons $\bm{S}^{\prime}$. They include:
@@
Fig. 3 shows examples of our simple heatmap augmentation techniques based on 2D skeletons. For position and orientation augmentations, we empirically use a multivariate Gaussian distribution with a covariance matrix so that the added noise is temporally smooth as in CASA [kwon2022context]. However, it does not lead to improved results and is thus not used in the above. Furthermore, the above augmentations generate correspondence labels between $\bm{S}$ and $\bm{S}^{\prime}$ (and similarly, between $\bm{H}$ and $\bm{H}^{\prime}$), which can be used for self-supervised training of our alignment network in the next section.
@@
Self-Attention. Given the original heatmap sequence $\bm{H}$ and the augmented sequence $\bm{H}^{\prime}$, we pass them separately through a video transformer, namely ViViT [arnab2021vivit], to perform self-attention both in the spatial and temporal domains, yielding effective spatiotemporal and contextual cues within each sequence. Below we discuss the processing of $\bm{H}$ only since the same is applied for $\bm{H}^{\prime}$. For each $\bm{h}_{i}$ in $\bm{H}$, we first break it into hwh*w non-overlapping patches, which are embedded via linear projection into hwh*w tokens. The tokens from $\bm{h}_{i}$ (with spatial positional encoding) are then fed to a spatial transformer, which learns interactions between tokens from the same temporal index (i.e., between joints in the same skeleton). Next, we employ global average pooling on the above tokens output by the spatial transformer to obtain the frame-level token representing $\bm{h}_{i}$. The frame-level tokens of $\bm{H}$ (with temporal positional encoding) are passed together to a temporal transformer, which learns interactions between tokens from different temporal indexes (i.e., between skeletons in the same sequence). Note that our approach performs self-attention both in the spatial and temporal domains, whereas CASA [kwon2022context] conducts self-attention in the temporal domain only. Fig. 4 illustrates this difference. In addition, our positional encodings are learned, whereas it is fixed in CASA [kwon2022context].
Cross-Attention. The above modules only focus on self-attention in the spatial and temporal domains within each sequence. Following [sun2021loftr, kwon2022context], we add a cross-attention module to learn interactions between $\bm{H}$ and $\bm{H}^{\prime}$ (i.e., between skeletons across the sequences), yielding contextual cues across the sequences. In particular, we encode the tokens output by the self-attention modules via an MLP layer and positional encoding before applying cross-attention.
Projection Head. We include a projection head, which is an MLP network with one hidden layer. As mentioned in [chen2020simple, kwon2022context], adding the projection head improves the generalization ability and yields effective features for downstream tasks.
Matching and Loss. We first compute the probability $\bm{\omega}_{ij}$ that the $i$-th frame in $\bm{H}$ is matched with the $j$-th frame in $\bm{H}^{\prime}$ using contrastive regression [dwibedi2019tcc, kwon2022context] as:
[…]